OpenStudy (anonymous):

The height of a triangle is 4 m less than its base. The area of the triangle is 48 m2. Find the length of the base. A. 7 m B. 8 m C. 11 m D. 12 m

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

when they said area of triangle is 48 do they mean the base?

OpenStudy (anonymous):

The area of a triangle equals 1/2 times base times height Let x represent the triangle's base 48 = 1/2 * x * (x-4) 96 = x^2 -4x x^2 -4x -96 = 0 factor the equation (x-12) * (x+8) = 0 x=12, x=-8 correct solution is the base equals 12 meters

OpenStudy (anonymous):

idk the bottom im guessing list im stupid and brain dead when it comes to math lol :)

OpenStudy (anonymous):

Let height be h Base = h + 4 Area = 48 1/2bh = 48 bh = 96 h(h + 4) = 96 h^2 + 4h - 96 = 0 (h + 12)(h - 8) = 0 h = 8 Base = 12m

OpenStudy (anonymous):

yay thank you both

OpenStudy (anonymous):

welcome and yall are slow i did that problem in under a minute

OpenStudy (anonymous):

area of a triangle=1/2*base* hieght.......so base =12

OpenStudy (anonymous):

k